# Semantic Self-Distillation for Language Model Uncertainty

### v1

@@ Abstract
Large language models present challenges for principled uncertainty quantification, in part due to their complexity and the diversity of their outputs. Semantic dispersion, or the variance in the meaning of sampled answers, has been proposed as a useful proxy for model uncertainty, but the associated computational cost prohibits its use in latency-critical applications. We show that sampled semantic distributions can be distilled into lightweight student models which estimate a prompt-conditioned uncertainty before the language model generates an answer token. The student model predicts a semantic distribution over possible answers; the entropy of this distribution provides an effective uncertainty signal for hallucination prediction, and the probability density allows candidate answers to be evaluated for reliability. On TriviaQA, our student models match or outperform finite-sample semantic dispersion for hallucination prediction and provide a strong signal for out-of-domain answer detection. We term this technique Semantic Self-Distillation (SSD), which we suggest provides a general framework for distilling predictive uncertainty in complex output spaces beyond language.
@@ 1 Introduction
Uncertainty quantification (UQ) can be used to assess the reliability of large language model (LLM) outputs (Xiong et al., 2024; Liu et al., 2025). As LLMs are rapidly deployed in high-stakes domains such as healthcare and law, it becomes increasingly important to have meaningful measures for the trustworthiness of model-generated content. Well-designed UQ methods can provide signals for detecting hallucinations - outputs that are misaligned with fact or user intention - and enable mitigation strategies such as abstention or optimal answer selection (Wen et al., 2025).
@@
LLMs are now frequently embedded in agentic frameworks, which chain multiple model invocations to solve complex, multi-step challenges (Yao et al., 2023; Xi et al., 2025). In these settings a single unreliable output can result in an error which compounds destructively over long-horizon tasks, resulting in complete failure of the system to achieve its goal (Kwa et al., 2025). Given the computational burden of agentic frameworks, practical uncertainty estimators must be low latency compared to the base language models employed.
@@
Semantic dispersion can be a useful proxy for uncertainty: when a language model generates multiple semantically distinct answers to the same prompt, the response is unlikely to be factual and correct. Measures based on semantic dispersion are effective for hallucination detection (Farquhar et al., 2024) but require generating many stochastic completions, rendering them prohibitively slow for latency-sensitive agentic applications.
@@
Probing techniques can address this latency issue by employing lightweight models to predict semantic dispersion or hallucination probability directly from internal model representations (Kadavath et al., 2022; Obeso et al., 2025; Han et al., 2025; Kossen et al., 2024). However, these methods typically collapse the complex notion of language model uncertainty into a single scalar metric, ignoring the topology of the underlying output distribution. This lossy reduction limits their application to simple risk classification, preventing the extraction of richer reliability signals.
@@
In this work, we propose Semantic Self-Distillation (SSD), a general UQ framework which bridges simple probing techniques and expensive sampling-based methods. Instead of regressing a single scalar statistic, SSD trains a small student model to predict a prompt-conditioned distribution over semantic answer embeddings. We model the student as a mixture density network (MDN) (Bishop, 1994), and approximate the teacher distribution using stochastic samples from the base language model. The resulting density acts as a general-purpose uncertainty object: its analytic entropy provides a pre-generation risk forecast, while its posterior likelihood enables post-generation answer verification and other distribution-level operations. Figure 1 provides an overview of the SSD training procedure and the inference-time applications investigated in this work.
On the TriviaQA dataset, SSD provides hallucination-risk signals that in many cases improve upon finite-sample empirical dispersion estimates. Beyond hallucination classification, the distilled density supports (i) context verification via likelihood-based out-of-domain detection and (ii) semantic consensus estimation from a single forward pass.
@@
  • We introduce Semantic Self-Distillation, a method for distilling a language model’s sampled semantic answer distribution into a lightweight conditional density estimator.
  • We evaluate SSD for pre-generation hallucination prediction, characterizing its cost-accuracy trade-offs relative to sampling-based semantic dispersion and scalar probes.
  • We demonstrate additional reliability primitives enabled by the predicted density, including post-generation verification via likelihood and single-pass semantic consensus estimation.
Figure 1: Overview of Semantic Self-Distillation (SSD). (A) Training: For each prompt, a hidden representation $\mathbf{h}$ conditions a student that distils sampled answer embeddings into a semantic mixture distribution. (B-D) Inference: The predicted density supports pre-generation hallucination prediction via entropy, post-generation context verification via likelihood, and semantic consensus estimation via the mixture mean. Plots illustrate the semantic embedding space, where points denote sampled answers and the mixture mean represents the model’s semantic consensus.
@@ 2.1 Problem Setup
Let $x_{1:T}$ be an input prompt and $y_{1:L}$ a completion generated by an autoregressive model. Let $\mathbf{h}=f_{\theta}(x_{1:T})\in\mathbb{R}^{d_{h}}$ denote a vector representation of the prompt extracted from the model’s internal states. We define a semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of the completed sequence via a mapping
@@
We train a student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ to predict the distribution over semantic embeddings $\mathbf{z}$ given only $\mathbf{h}$. The entropy of $q_{\phi}(\cdot\mid\mathbf{h})$ provides an uncertainty signal for hallucination prediction.
2.2 Prompt and Answer Representations
Given a user prompt $x$ of length $T$, we extract hidden states directly from the base language model. Let $\mathbf{h}_{t}^{(\ell)}\in\mathbb{R}^{d}$ denote the hidden state at layer $\ell$ for token position $t$. We utilize the representation of the final token of the prompt extracted from a single specific layer $\ell^{*}$:
@@ 2.2 Prompt and Answer Representations
The layer index $\ell^{*}$ is treated as a hyperparameter. In practice, we estimate the optimal $\ell^{*}$ by identifying the layer most predictive of the model’s semantic uncertainty, as detailed in Section 3.
For the answer, we apply a specialist embedding model to the generated sequence to obtain the semantic representation (see Section 3 for model details). To improve training efficiency and focus on the principal axes of semantic variation, we apply Principal Component Analysis (PCA) to the target embeddings $\mathbf{z}$. We fit the PCA transformation on the flattened set of stochastic training completions.
2.3 Student Distribution
We model the student as an MDN (Bishop, 1994). Conditioned on the prompt representation $\mathbf{h}\in\mathbb{R}^{d_{h}}$, the student predicts a $K$-component Gaussian mixture over the semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$:
@@ 2.3 Student Distribution
where an MLP with parameters $\phi$ maps $\mathbf{h}$ to mixture logits $\bm{\alpha}(\mathbf{h})\in\mathbb{R}^{K}$, component means $\bm{\mu}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$, and (log-)scales $\bm{\sigma}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$. We train by maximizing the conditional log-likelihood of teacher samples. For data efficiency, we use diagonal covariances.
2.4 Teacher Construction and Training
@@
For each training prompt $x$, we extract the representation $\mathbf{h}$ (using the selected layer $\ell^{*}$) and generate $S$ stochastic completions $y^{(1)},\dots,y^{(S)}$. We compute semantic embeddings $\mathbf{z}^{(s)}=\Phi(y^{(s)})$ for $s{=}1,\dots,S$, and fit the student parameters $\phi$ by conditional maximum likelihood:
$\max_{\bm{\phi}}\ \sum_{\mathbf{x}}\ \frac{1}{S}\sum_{s=1}^{S}\log q_{\bm{\phi}}\!\big(\mathbf{z}^{(s)}\mid\mathbf{h}\big).$
2.5 Quantifying Dispersion
Given a new test prompt $x_{1:T}$, we compute its representation $\mathbf{h}=f_{\theta}(x_{1:T})$ in a single forward pass of the base model. The student then predicts the parameters of a $K$-component diagonal-covariance Gaussian mixture over semantic embeddings. We measure dispersion using the order-2 Rényi entropy (Rényi, 1961) of the predictive density:
$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z},$
Intuitively, the integral $\int q(\mathbf{z})^{2}d\mathbf{z}$ is the collision probability of the density: it is larger when mass is concentrated, and smaller when the distribution is spread out (Principe, 2010). Higher $\mathsf{H}_{2}$ corresponds to greater semantic dispersion.
We use Rényi-2 because it yields an exact closed form for Gaussian mixtures (Wang et al., 2009), enabling a sampling-free, analytic dispersion score at inference time. By contrast, the Shannon differential entropy (Shannon, 1948) does not, in general, admit a closed-form expression for Gaussian mixture models; it is typically approximated using Monte Carlo estimation or via analytic bounds (Huber et al., 2008; Dahlke and Pacheco, 2023). Using Rényi-2 therefore preserves SSD’s goal of avoiding additional sampling or expensive numerical estimation at deployment. We provide the closed form expression and derivation in Appendix Section A.
2.6 Using the Predicted Semantic Distribution
@@
At inference time, SSD produces a prompt-conditioned semantic density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ over answer embeddings, which we treat as a general-purpose uncertainty object with three primitive operations:
(i) Semantic Dispersion Prediction
@@
We use the Rényi-2 entropy $\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)$ to provide a prompt-level uncertainty estimate in a single forward pass.
(ii) Context Verification
@@
Given any candidate response $\mathbf{y}^{\prime}$, we embed it as $\mathbf{z}^{\prime}=\Phi(\mathbf{y}^{\prime})$ and score its compatibility with the prompt via the posterior log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$. We treat answers with low likelihood as out-of-domain relative to the prompt context.
(iii) Latent Sampling
@@
Because $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ is an explicit mixture distribution, we can compute summary statistics such as the mixture mean $\mathbb{E}_{q_{\phi}}[\mathbf{z}]=\sum_{k}\pi_{k}(\mathbf{h})\,\bm{\mu}_{k}(\mathbf{h})$ and, when desired, draw latent samples $\tilde{\mathbf{z}}\sim q_{\phi}(\cdot\mid\mathbf{h})$.

### v2

@@ Abstract
Large language models present challenges for principled uncertainty quantification, in part due to their complexity and the diversity of their outputs. Semantic dispersion, or the variance in the meaning of sampled answers, has been proposed as a useful proxy for model uncertainty, but the associated computational cost prohibits its use in latency-critical applications. We show that sampled semantic distributions can be distilled into lightweight student models which estimate a prompt-conditioned density before the language model generates an answer token. The student model predicts a semantic distribution over possible answers; the entropy of this distribution provides a prompt-level uncertainty signal, and the probability density allows answer-level reliability evaluation. Across experiments on TriviaQA and MMLU, we find our student models perform competitively relative to sampling-based semantic dispersion baselines on a hallucination prediction task, whilst offering additional uncertainty primitives for out-of-domain detection and multiple-choice answer selection. We term this technique Semantic Self-Distillation (SSD), which can serve as a general framework for distilling predictive uncertainty in complex output spaces beyond language.
@@ 1 Introduction
Figure 1: Overview of Semantic Self-Distillation (SSD). (A) Training: For each prompt, a hidden representation $\mathbf{h}$ conditions a student that distils sampled answer embeddings into a semantic mixture distribution. (B-D) Inference: The predicted density supports pre-generation hallucination prediction via entropy, post-generation context verification via likelihood, and multiple-choice answer selection. Plots illustrate the semantic embedding space, where points denote sampled answers.
@@
Uncertainty quantification (UQ) can be used to assess the reliability of large language model (LLM) outputs [xiong2024efficient, liullmuncertainty2025]. As LLMs are rapidly deployed in high-stakes domains such as healthcare and law, it becomes increasingly important to have meaningful measures for the trustworthiness of model-generated content. Well-designed UQ methods can provide signals for detecting errors like hallucinations - outputs that are misaligned with fact or user intention - and enable mitigation strategies such as abstention, information retrieval, or answer selection [wen-etal-2025-know].
@@
LLMs are now frequently embedded in agentic frameworks, which chain multiple model invocations to solve complex, multi-step tasks [yao2023react, xi2025rise]. In these settings a single unreliable output can introduce an error that compounds destructively over long-horizon tasks, potentially resulting in complete failure of the system to achieve its goal [kwa2025measuring]. Given the computational burden of such agentic frameworks, practical uncertainty estimators must be low latency compared to the base language models.
@@
Semantic dispersion has emerged as a useful proxy for uncertainty: when a language model generates semantically diverse answers to the same prompt, the response is unlikely to be factual and correct. Sampling-based measures of semantic dispersion are effective for hallucination detection [farquhar_detecting_2024], but require generating multiple stochastic completions, rendering them prohibitively expensive in latency-sensitive applications.
@@
Probing techniques can address this latency issue by training lightweight models to predict semantic dispersion or hallucination probability directly from internal model representations [kadavath2022languagemodelsmostlyknow, obeso2025realtimedetectionhallucinatedentities, han-etal-2025-simple, kossen2024semanticentropyprobesrobust]. However, these methods typically collapse the complex notion of language model uncertainty into a single scalar metric, ignoring the topology of the underlying output distribution. This information reduction limits their application to simple risk classification, preventing the extraction of richer reliability signals.
@@
In this work, we propose Semantic Self-Distillation (SSD), a general UQ framework that bridges lightweight probing techniques and expensive sampling-based methods. Instead of regressing a scalar uncertainty statistic, SSD trains a small student model to predict a prompt-conditioned distribution over semantic answer embeddings. We model the student as a mixture density network (MDN) [bishop1994mixture], and approximate the teacher distribution using stochastic samples from the base language model. The predicted density serves as a general-purpose uncertainty object: its analytic entropy provides a pre-generation risk forecast, whilst estimated likelihoods enable post-generation answer verification. Figure 1 provides an overview of the SSD training procedure and the inference-time applications investigated in this work.
We evaluate SSD on a hallucination prediction task with TriviaQA and MMLU across multiple LLM families, demonstrating competitive, and in some cases even superior, performance relative to sampling-based semantic dispersion baselines. While specialized scalar probes can be strong pre-generation classifiers, SSD uniquely retains distributional structure in a single forward pass. This enables additional post-generation capabilities unavailable to standard single-pass methods, which we demonstrate by evaluating SSD on likelihood-based out-of-domain answer detection and multiple-choice answer selection.
@@
  • We introduce Semantic Self-Distillation (SSD), a method for distilling a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned density estimator.
  • We show that SSD provides competitive pre-generation hallucination prediction performance across multiple model families, while matching the inference latency of single-pass probing methods.
  • We demonstrate that modelling the full semantic distribution unlocks new reliability primitives beyond scalar uncertainty scoring, including post-generation answer verification and distribution-based answer selection.
@@ 2 Methods
In this section, we describe how SSD distils a language model’s sampled semantic output distribution into a prompt-conditioned density estimator, rather than a scalar uncertainty proxy. We first define the representations used for prompts and answers, specify the MDN student model, and detail the distillation objective used for training. We then show how the resulting predictive density can be used to extract both pre-generation, prompt-level dispersion estimates, and post-generation, answer-level likelihood scores, all in a single forward pass at inference time.
@@ 2.1 Problem Setup
Let $x_{1:T}$ be an input prompt and $y_{1:L}$ a completion generated by a decoding model. Let $\mathbf{h}=f_{\theta}(x_{1:T})\in\mathbb{R}^{d_{h}}$ denote a vector representation of the prompt extracted from the model’s internal states. We define a semantic representation $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of the completed sequence via a mapping $\Phi$, which in practice is an off-the-shelf sequence embedding model:
@@
We train a student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ to predict the distribution over semantic embeddings $\mathbf{z}$ given only $\mathbf{h}$.
2.1.1 Prompt and Answer Representations
Given a user prompt $x_{1:T}$ of length $T$, we extract hidden states directly from the base language model. Let $\mathbf{h}_{t}^{(\ell)}\in\mathbb{R}^{d}$ denote the hidden state at layer $\ell$ for token position $t$. We utilize the representation of the final token of the prompt extracted from a single specific layer $\ell^{*}$:
@@ 2.1.1 Prompt and Answer Representations
The layer index $\ell^{*}$ is treated as a hyperparameter. In practice, we estimate the optimal $\ell^{*}$ by identifying the layer most predictive of the model’s semantic uncertainty, as explained in Appendix D.1.
For the answer, we apply a specialist embedding model $\Phi$ to the generated sequence $y_{1:L}$ to obtain the semantic representation, $\mathbf{z}=\Phi(y_{1:L})$, with full model details provided in Section 4.1. To improve training efficiency and focus on the principal axes of semantic variation, we also apply Principal Component Analysis (PCA) to reduce the dimensionality of the target embeddings $\mathbf{z}$.
2.1.2 Student Model
We define the student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ using an MDN [bishop1994mixture]. Conditioned on the prompt representation $\mathbf{h}\in\mathbb{R}^{d_{h}}$, the student predicts a $K$-component Gaussian mixture over the semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$:
@@ 2.1.2 Student Model
where an MLP with parameters $\phi$ maps $\mathbf{h}$ to mixture logits $\bm{\alpha}(\mathbf{h})\in\mathbb{R}^{K}$, component means $\bm{\mu}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$, and scales $\bm{\sigma}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$. For data efficiency, we use diagonal covariance matrices.
@@
For each training prompt $x$, we extract the representation $\mathbf{h}$ (using the selected layer $\ell^{*}$) and generate $S$ stochastic completions $y^{(1)},\dots,y^{(S)}$ using the base language model. We compute semantic embeddings $\mathbf{z}^{(s)}=\Phi(y^{(s)})$ for $s{=}1,\dots,S$, and fit the student parameters $\phi$ by conditional maximum likelihood:
$\max_{\bm{\phi}}\ \sum_{x}\ \frac{1}{S}\sum_{s=1}^{S}\log q_{\bm{\phi}}\!\big(\mathbf{z}^{(s)}\mid\mathbf{h}(x)\big).$
2.2 Extracting Uncertainty Signals
@@
At inference time, SSD produces a prompt-conditioned semantic density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ over answer embeddings, which we treat as a general-purpose uncertainty object. In this work, we focus on two primary operations derived from this object: prompt-level dispersion and answer-level likelihood. We also highlight that the explicit density enables additional distributional operations such as sampling.
@@
(i) Prompt evaluation. We use the entropy of the predicted semantic distribution $q_{\phi}(\cdot\mid\mathbf{h})$ to provide a prompt-level uncertainty estimate in a single forward pass.
@@
(ii) Answer evaluation. Given any candidate response $\mathbf{y}^{\prime}$, we embed it as $\mathbf{z}^{\prime}=\Phi(\mathbf{y}^{\prime})$ and score its compatibility with the prompt via the posterior log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$. We may treat answers with low likelihood as out-of-domain relative to the prompt context.
@@
(iii) Latent sampling. Because $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ is an explicit mixture distribution, we can compute summary statistics such as the mixture mean $\mathbb{E}_{q_{\phi}}[\mathbf{z}]=\sum_{k}\pi_{k}(\mathbf{h})\,\bm{\mu}_{k}(\mathbf{h})$ and, when desired, draw latent samples $\tilde{\mathbf{z}}\sim q_{\phi}(\cdot\mid\mathbf{h})$.
We investigate operations (i) and (ii) in Sections 4.2 and 4.3, respectively. Operation (iii) illustrates a structural consequence of modelling the full predictive density and is discussed in Appendix B.2.
2.2.1 Quantifying Dispersion
We measure dispersion using the order-2 Rényi entropy [renyi1961measures] of the predictive density $q_{\phi}(\cdot\mid\mathbf{h})$:
$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z}.$
Intuitively, the integral $\int q(\mathbf{z})^{2}d\mathbf{z}$ is the collision probability of the density: it is larger when mass is concentrated, and smaller when the distribution is spread out [principe2010information]. Higher $\mathsf{H}_{2}$ corresponds to greater semantic dispersion.
We use Rényi-2 because it yields an exact closed form for Gaussian mixtures [wang2009closed], enabling a sampling-free, analytic dispersion score at inference time. By contrast, the Shannon differential entropy [shannon1948mathematical] does not, in general, admit a closed-form expression for Gaussian mixture models; it is typically approximated using Monte Carlo estimation or via analytic bounds [huber2008entropy, dahlke2023closedform]. Using Rényi-2 therefore preserves SSD’s goal of avoiding additional sampling or expensive numerical estimation at deployment. We provide the closed form expression and derivation in Appendix A.
2.3 Extension to Generic Sequence Domains
Although we focus on autoregressive language models in this work, the SSD construction is agnostic to modality. In language tasks, $\mathbf{z}=\Phi(y_{1:L})$ captures the semantic content of a generated answer. More generally, $\mathbf{z}$ may represent any task-relevant latent characterization of a predicted sequence $y_{1:L}$, rather than its individual tokens or events.
Under stochastic generation, a sequence model induces a predictive distribution over such latent representations. SSD distils this sampled distribution into a lightweight conditional density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$, where $\mathbf{h}$ encodes the input sequence. The resulting density provides a sampling-free approximation to the model’s belief over high-level outcomes. We provide further discussion of potential application domains in Section 5.
3 Related Work
Uncertainty estimation in language models.
xiong2024efficient categorize UQ methods for LLMs into three broad categories: single-sample, multi-sample, and probing-based. Single-sample methods are training-free and include output perplexity, negative sequence probability, and mean token entropy [malinin2021uncertainty]. Probing-based methods similarly use a single sample at inference time but train a predictive model on internal LLM activations to predict an uncertainty-relevant quantity, such as the probability of correctness [kadavath2022languagemodelsmostlyknow] or semantic dispersion [kossen2024semanticentropyprobesrobust]. Multi-sample methods often rely on notions of semantic dispersion, either explicitly, such as Semantic Entropy [farquhar_detecting_2024] and Number of Clusters [lin2024generating], or implicitly, as with P(true) [kadavath2022languagemodelsmostlyknow]. We emphasize that these approaches typically model a proxy for language model uncertainty, geared for application in a particular task. There is limited consensus on how to decompose classical notions of aleatoric and epistemic uncertainty in LLMs, and the resulting uncertainty signals are often only modestly correlated [hou2024decomposing, xiong2024efficient].
Distillation in language models.
Distillation typically refers to training smaller language models on the logit or token outputs of more capable models, with notable success in reasoning domains [guo2025deepseek]. Such approaches transfer predictive behaviour but do not explicitly target calibrated uncertainty or structured output distributions beyond the token level. In contrast, we focus on sequence-level distributions of model outputs and investigate whether these can be elicited without autoregressive sampling. piskorz2026eliciting study a related but narrower phenomenon concerning language model predicted distributions over numerical targets, whereas we consider broader semantic targets.
Knowledge distillation for uncertainty estimation.
A related line of work seeks to distil ensemble-based uncertainty estimates into a single model [Malinin2020Ensemble, fathullah2022self, landgraf2024dudes]. Ensembles are widely regarded as producing high-quality uncertainty estimates, but incur substantial computational cost at inference time. Recent approaches therefore train a student model to approximate ensemble predictive distributions or uncertainty signals, aiming to recover ensemble-quality uncertainty in a single forward pass, typically by matching logits, predictive variances, or other distributional summaries [fathullah2023logit, park2025knowledge].
In contrast, our setting does not assume access to an external ensemble and does not aim to replicate variance estimates or token-level predictive distributions. Instead, we distil a language model’s semantic sequence-level output distribution induced by stochastic sampling, learning a prompt-conditioned density over answer embeddings. The most closely related approaches in the LLM literature are probing-based uncertainty estimators [kadavath2022languagemodelsmostlyknow, kossen2024semanticentropyprobesrobust], which however compress uncertainty into a single scalar statistic. By contrast, we predict a full sequence-level distribution in a single forward pass, enabling downstream operations beyond scalar risk scoring, including likelihood-based post-generation verification and multiple-choice answer selection. Rather than estimating a task-specific uncertainty heuristic, we approximate the underlying semantic output distribution itself.
Figure 2: Qualitative analysis of the SSD density. (A) A correct answer yields minimal SSD entropy (0.1st percentile), while a hallucination yields high uncertainty (99.7th), aligning with sampling-based Teacher Dispersion (TD) and Semantic Entropy (SE). Percentiles denote rank within the TriviaQA test set (N=1, [^0] N{=}1{,}000). (B) A precise prompt (P1) yields significantly lower entropy than an underspecified one (P2). (C) The SSD likelihood enables context-aware scoring: given identical candidate sets, the student assigns the highest likelihood to the semantically appropriate answer for each prompt.